# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v1

@@ IV.2.1 Single-cluster approximation for $J_{\text{MF}}$ and $J_{0}$
Figure 6: Justification for the single cluster approximation: (a) illustrates the one cluster regime, where distance $d$ is the part of the cluster that breaks off. When $\varepsilon\gg k^{-},k^{+}$, the break off travels around the circle before a new unpausing event. At low densities, a single paused particle can block the cluster, making $d$ cover the entire cluster, which upon an unpausing event, leads to the system with no paused particles. Thus, the system behaves like a standard TASEP. (b) shows the probability distribution of the number of clusters for $L=250$, $k^{-}=10^{-3}$, $k^{+}=10^{-6}$, $\rho=0.33$. The distribution is bimodal: the right side is Gaussian, confirming the standard TASEP phase, while the left side shows a large peak indicating a single cluster of particles.
@@
$d=\frac{k_{-}}{k_{+}}\left(1-\left(\frac{k_{-}}{k_{-}+k_{+}}\right)^{\rho L}\right)\,,$ (13)
which reduces to $k_{-}/k_{+}$ for clusters of infinite length. The value of $d$ can then be plugged into Eq. (12) to evaluate $J_{c}$.
@@
We are interested in the current from time 00, that is when the fully packed cluster (density $1$) starts to dissolve due to the absence of paused particles, to the expected time required for a new pause, which on average will occur after a time $\tau=\frac{1}{\rho Lk_{+}}$ that corresponds to the inverse of the rate of pausing any of the $\rho L$ particles. By solving the Burgers’ equation with the method of characteristics, we obtain $\rho(x,t)$ and we compute the integral Eq. (14). We obtain:
@@
evaluated at $\tau=\frac{1}{\rho Lk_{+}}$.
We now have all components of Eq. (11), which can thus be tested against numerical results. In Fig. 5 we plot the result of Eq. (11), using $J_{0}$ from Eq. (16) and $J_{c}$ from the single-cluster approximation Eq. (12), compared to both simulations and the mean-field theory. Our framework not only correctly describes the regimes when finite size effects are relevant but also captures them quantitatively.
@@ IV.2.2 Comparison with previous approaches
Figure 7: Relative error between the maximal value of the simulated current and the predicted value at the corresponding particle density. Panel (a) shows a heatmap for the error in the single-cluster solution given by Eq.( 12). Panels (b)-(c) show the ddTASEP and pTASEP using the mean-field approach, respectively. In panel (d) we emphasise how the parameter region where the approach fails shrinks with increasing the system size. We notice how previous approaches cannot capture the regime of long ($k_{-}\ll\varepsilon$) and rare ($k_{+}\ll\varepsilon$) pauses. For panel (a)-(c), $L=250$.
Our single-cluster approximation succeeds in predicting the system’s behaviour in cases of strong finite size effects, as shown in Fig. 5. We now systematically asses the quality of the single-cluster approximation in the $\{k_{-},k_{+}\}$ parameter space by comparing the particle current predicted by our approach to that obtained from simulations. Figure 7(a) shows a ‘heatmap’ of the relative error between $J_{max}$ obtained with simulations and with Eq. (11). For comparison we plot the same quantity for previously proposed mean-field approaches: the ddTASEP as studied in [11] -Fig. 7(b)- and $J_{\text{MF}}$ [9] for the pTASEP -Fig. 7(c). The models from the literature fail to capture finite-size effects occurring in the bottom left region, compared to the single cluster model introduced in the previous section. The abruptly increasing error for large $k_{-}$ and $k_{+}$, for the single-cluster model, is due to the failure of the single-cluster assumption in regimes where the timescale separation between those rates and $\varepsilon$ is less clear. In this regime, the single-cluster approximation obviously fails, being in a sense complementary to previous mean-field descriptions of the pTASEP and ddTASEP that cannot capture finite-size effects.
To further emphasise the fact that former approaches break down because of finite-size effects, in Fig. 7(d) we show the contour lines at $50\%$ relative error of $J_{\text{MF}}$ for different values of $L$, pointing out how the region with the largest disagreement (small $k_{-}$, $k_{+}$ with respect to $\varepsilon$) shrinks as the system size increases.
@@ IV.2.3 Extending the mean field in a multi-cluster regime
The single cluster approximation fails for increasing $k_{-}$, $k_{+}$, and $L$. On the one hand, when $k_{-}$ increases any cluster will break up before it can collect all the particles, leading to the emergence of multiple clusters. On the other hand, an increase in $k_{+}$ results in more frequent pauses, resulting in the inability for all the particles to coagulate into a single cluster during the time $1/k_{-}$. The system size ($L$) further exacerbates these issues by extending the time required for all the particles to form a single cluster. Consequently, both terms in Eq (12) fail due to the emergence of multiple clusters.
Figure 8: Relative error (at the maximum) between simulations and theory, Eq. (17). The extended mean-field equation works reasonably well for a large range of $k_{-}$ and $k_{+}$. In this plot $L=250$.
To address these limitations, we propose an effective mean-field approach. Starting from Eq. (12), we replace the single-block current $J_{c}$ with the mean-field current $J_{\text{MF}}$ given by Eq. (4) and the current term with no paused particles $J_{0}$ with the standard TASEP current given by Eq. (1). The resulting expression for the current $J$ is:
@@
It is important to note that, while this equation does not explicitly depend on the system size, the weight $P_{0}$ does. As seen in Figure 8, Eq. (17) works well over various orders of magnitude for $k_{-}$ and $k_{+}$, indicating the robustness of our approach.
@@ IV.3 Open boundary conditions
Treating open boundaries is considerably more complex than the periodic case, as the number of particles on the lattice is no longer constant over time. As a consequence, we can neither use the number of particles to compute probabilities nor assume that the pausing dynamics has reached its steady state. Within the parameter range described in Section IV.1 one would expect the single-blockage phenomenology explored from the periodic boundary case translates to the open boundary case. The lattice can then be found in either of two states: a ‘clogged’ regime (bearing similarities with the clustered regime in periodic boundary conditions), where the lattice is completely filled up to the site of the first paused particle, and a pause-free state, where the system behaves as a standard TASEP. Because of the time separation between $k_{-}$, $k_{+}$ and $\varepsilon$, we can assume that particles quasi-instantaneously fill the lattice up to the position of the first particle becoming paused. Similarly, and for the same reason, particles downstream from the rightmost paused particle would quasi-instantaneously exit the lattice. Reproducing the reasoning followed to obtain Eq. (11), we therefore introduce a two term approach:
$J=\frac{\tau_{0}}{\tau_{c}+\tau_{0}}J_{0}+\frac{\tau_{c}}{\tau_{c}+\tau_{0}}J_{c}\,,$ (18)
@@
where $\tau_{0}$ and $\tau_{c}$ are the average duration of the pause-free state and of the clogged state, respectively. The pause-free phase behaves as a standard TASEP. Thus, $J_{0}$ is given by the standard TASEP current Eq.(1).
@@
We denote by $x$ the position at which, starting from the pause-free regime, a particle transitions to the paused state. The position $x$ (in units of lattice sites) also corresponds to the total number of particles constituting the initial cluster. Considering that in the standard TASEP the particle density is homogeneous, and since all particles have the same probability to transition towards the paused state, we have $x\right>=L/2$ on average. Thus the total current of the clogged regime can be computed as the number of particles in the initial cluster divided by the typical time $\tau_{c}$ of its duration, $J_{c}=L/(2\tau_{c})$. Equation (18) then becomes:
$J=\frac{\tau_{0}}{\tau_{0}+\tau_{c}}J_{0}+\frac{1}{\tau_{0}+\tau_{c}}\frac{L}{2}.$
@@
The first timescale $\tau_{0}$ can be computed by considering that the rate of switching from the pause-free state is proportional to the total number of particles times $k_{+}$, that is
$\tau_{0}=\frac{1}{\rho Lk_{+}}=\frac{\varepsilon}{\alpha Lk_{+}}\,,$
@@
where in the last equivalence we consider a lattice in the LD phase ($\rho=\alpha/\varepsilon$). We then compute the mean lifespan $\tau_{c}$ of the clogged regime following a numerical approach. This timescale can be smaller or comparable to the time necessary to reach an equilibrium distribution of the paused particle fraction $f_{p}$ given in Eq. (3), and for this reason writing a closed form for $\tau_{c}$ is cumbersome. Thus, after the first pausing event leading to the formation of the extensive cluster, other particles can become paused, slowing down the restoration of the pause-free regime. Eventually, the clogged regime will disappear when all particles will successively re-become active and exit the lattice together with the particles that were blocked behind them. The silent assumption behind our reasoning is that active particles which are not trapped in the cluster will exit the lattice without becoming paused again.
@@
To develop our numerical method we first compute the time dependence of $f_{p}$ starting from a situation in which all particles are active:
$f_{p}(t)=\frac{k_{+}}{k_{+}+k_{-}}\left(1-e^{-(k_{+}+k_{-})t}\right)\,,$ (19)
@@
which asymptotically reaches the steady-state probability $f_{p}=\frac{k_{+}}{k_{+}+k_{-}}$. At a given time $t$, the mean number $d_{t}$ of particles ‘trapped’ between the first and the second rightmost paused particles can be computed as the mean number of failures in a sequence of independent and identically distributed Bernoulli trials before another particle is found paused (success). This is computed as the average of a negative binomial distribution with rate of success $f_{p}(t)$ [Eq. (19], giving
@@
The average lifetime $\tau(x)$ of a paused particle situated at $x$ is $t_{-}:=1/k_{-}$, and the expected number of particles between two paused particles is given by Eq. (20) computed at $t=t_{-}$. We now have all the elements to numerically compute the typical time to free all particles blocked in the clogged regime. The number of particles freed after the leading particle becomes active is given by $d_{1}+1$, where $d_{1}$ is the number of particles in between the first and the second paused particles when the leading particle becomes active at time $t_{1}:=1\times t_{-}$. If $d_{1}+1$ is greater than $x$, we stop and estimate $\tau(x)=t_{-}$. Otherwise, we compute the next expected distances $d_{i}$ evaluating Eq.(20) at $t_{i}=i\times t_{-}$ until $\sum_{i=1}^{j}(d_{i}+1)>x$. Then, we approximate $\tau(x)$ as $j\times t_{-}$.
Averaging $\tau(x)$ over all possible cluster lengths finally gives us the $\tau_{c}$ to be used in Eq. (18). While this estimate is indeed numerical it has to be considered that it is still considerably faster than directly simulating the system, and its success proves the predominance of finite size effects and the validity of our approximation.
Figure 9: Compares the current as a function of $k_{+}$ for different system sizes. Here $k_{-}=10^{-3},\alpha=0.1,$ and $\beta=1$. The solution given by Eq. 18 works well for small system sizes and becomes worse for larger lattices.
Our theory is in good agreement with the simulations, particularly for small lattices. In particular, in Fig. 9 we show how our two-state approximation Eq. (18) correctly predicts the particle current $J$ even for small systems ($L=50$) and when the pausing rate $k_{+}$ is varied over several orders of magnitudes.
@@ V Discussions
Intriguingly, in both periodic and open boundary conditions (see Figs. 3-4), this mean-field approximation fails when $k_{\pm}/\varepsilon\ll 1$. We provide a few estimates of rates involved in transcription and translation, and point out that biologically relevant regimes are close to, or fall into, this region of poor agreement between theory and simulations. Corroborated by a systematic exploration of the parameter space (Figs. 5-7), we conclude that severe finite-size effects cause the inadequacy of existing mean-field approaches when $k_{\pm}/\varepsilon\ll 1$. In contrast to the standard TASEP, where the mean-field theory performs well even with relatively small lattices (of the order of $100$ sites), and where the agreement with simulations is independent of hopping and boundary rates, the pTASEP exhibits finite-size effects that depend on parameter values. While the thermodynamic limit remains a reasonable approximation for finite systems, the required system size to achieve this condition is closely tied to the separation of timescales between pausing/unpausing dynamics and particle hopping.
We develop a single-cluster approximation (Section IV.2.1) to better understand the nature of the interplay between this timescale separation and finite-size effects. As the size dependence is related to the disappearance of paused particles in the lattice, our method, which considers the lattice as being in two states (with or without paused particles), captures the relation between current and lattice size. To develop this, in the regime $k_{\pm}/\varepsilon\ll 1$ we assume the extreme cases of all particles regrouping in a single cluster, or relaxing to the standard TASEP dynamics when there are no paused particles. This single-cluster approximation reproduces well the simulated current-density relationships when the hypotheses are met (Figs. 5-7). However, the relaxation process (necessary to achieve the steady-state current) being fast compared to the pausing/unpausing kinetics (see Fig. 11), we can generalise our approach to the multi-cluster regime, for which we approximating the current $J_{0}$ in the absence of paused particles by that of the standard TASEP Eq. (1).
@@
Finally, we present the extension of the single-cluster approximation to the open boundary conditions, showing how our two-state hypothesis can be extended to this situation. For instance, our model describes well how the current depends on the pausing rate $k_{+}$ across different orders of magnitudes.

### v2

@@ IV.2.1 Single-cluster approximation for $J_{\text{p}}$ and $J_{0}$
Figure 6: Justification for the single cluster approximation: (a) illustrates the single cluster regime, where distance $d$ is the part of the cluster that breaks off. When $\varepsilon\gg k_{u},k_{p}$, the break off travels around the circle before a new unpausing event. At low densities, a single paused particle can block the cluster, making $d$ cover the entire cluster, which upon an unpausing event, leads to the system with no paused particles. Thus, the system behaves like a standard TASEP. (b) shows the probability distribution of the number of clusters for $L=250$, $k_{u}=10^{-3}$, $k_{p}=10^{-6}$, $\rho=0.33$. The distribution is bimodal: the right side is Gaussian, confirming the standard TASEP phase (unpaused state), while the left side shows a large peak indicating a single cluster of particles.
@@
$d=\frac{k_{u}}{k_{p}}\left(1-\left(\frac{k_{u}}{k_{u}+k_{p}}\right)^{\rho L}\right)\,,$ (13)
which reduces to $k_{u}/k_{p}$ for clusters of infinite length. The value of $d$ can then be plugged into Eq. (12) to evaluate $J_{p}$.
@@
We are interested in the current from time 0, that is when the fully packed cluster (density 111) starts to dissolve due to the absence of paused particles, to the expected time required for a new pause, which on average will occur after a time $\tau=\frac{1}{\rho Lk_{p}}$ that corresponds to the inverse of the rate of pausing any of the $\rho L$ particles. By solving the Burgers’ equation with the method of characteristics, we obtain $\rho(x,t)$ and we compute the integral Eq. (14). We obtain:
@@
evaluated at $\tau=\frac{1}{\rho Lk_{p}}$.
We now have all components of Eq. (11), which can thus be tested against numerical results. In Fig. 5 we plot the result of Eq. (11), using $J_{0}$ from Eq. (16) and $J_{p}$ from the single-cluster approximation Eq. (12), compared to both simulations and the mean-field theory. Our framework not only correctly describes the regimes when finite size effects are relevant but also captures them quantitatively.
@@ IV.2.2 Comparison with previous approaches
Figure 7: Relative error between the maximal value of the simulated current and the predicted value at the corresponding particle density, computed as $|J-J_{\text{sim}}|/|J_{\text{sim}}|$, where $J$ represents the analytical solution for the current and $J_{\text{sim}}$ is the simulated current. Panel (a) shows a heatmap for the error in the single-cluster solution given by Eq.( 12). Panels (b)-(c) show the ddTASEP and pTASEP using the mean-field approach, respectively. In panel (d) we emphasise how the parameter region where the approach fails shrinks with increasing the system size. We notice how previous approaches cannot capture the regime of long ($k_{u}\ll\varepsilon$) and rare ($k_{p}\ll\varepsilon$) pauses. For panel (a)-(c), $L=250$.
Our single-cluster approximation succeeds in predicting the system’s behaviour in cases of strong finite size effects, as shown in Fig. 5. We now systematically asses the quality of the single-cluster approximation in the $\{k_{u},k_{p}\}$ parameter space by comparing the particle current predicted by our approach to that obtained from simulations. Figure 7(a) shows a ‘heatmap’ of the relative error between $J_{max}$ obtained with simulations and with Eq. (11). For comparison we plot the same quantity for previously proposed mean-field approaches: the ddTASEP as studied in [11] -Fig. 7(b)- and $J_{\text{MF}}$ [9] for the pTASEP -Fig. 7(c). The models from the literature fail to capture finite-size effects occurring in the bottom left region, compared to the single cluster model introduced in the previous section. The abruptly increasing error for large $k_{u}$ and $k_{p}$, for the single-cluster model, is due to the failure of the single-cluster assumption in regimes where the timescale separation between those rates and $\varepsilon$ is less clear. In this regime, the single-cluster approximation obviously fails, being in a sense complementary to previous mean-field descriptions of the pTASEP and ddTASEP that cannot capture finite-size effects.
[…]
@@ IV.2.3 Extending the mean field in a multi-cluster regime
The single cluster approximation fails for increasing $k_{u}$, $k_{p}$, and $L$. On the one hand, when $k_{u}$ increases any cluster will break up before it can collect all the particles, leading to the emergence of multiple clusters. On the other hand, an increase in $k_{p}$ results in more frequent pauses, resulting in the inability for all the particles to coagulate into a single cluster during the time $1/k_{u}$. The system size ($L$) further exacerbates these issues by extending the time required for all the particles to form a single cluster. Consequently, both terms in Eq (12) fail due to the emergence of multiple clusters.
Figure 8: Relative error (at the maximum) between simulations and theory, Eq. (17). The extended mean-field equation works reasonably well for a large range of $k_{u}$ and $k_{p}$. In this plot $L=250$.
To address these limitations, we propose an effective mean-field approach. Starting from Eq. (12), we replace the single-block current $J_{p}$ with the mean-field current $J_{\text{MF}}$ given by Eq. (4) and the current term with no paused particles $J_{0}$ with the standard TASEP current given by Eq. (1). The resulting expression for the current $J$ is:
@@
It is important to note that, while this equation does not explicitly depend on the system size, the weight $P_{0}$ does. As seen in Figure 8, Eq. (17) works well over various orders of magnitude for $k_{u}$ and $k_{p}$, indicating the robustness of our approach.
@@ IV.3 Open boundary conditions
Treating open boundaries is considerably more complex than the periodic case, as the number of particles on the lattice is no longer constant over time. As a consequence, we can neither use the number of particles to compute probabilities nor assume that the pausing dynamics has reached its steady state. Within the parameter range described in Section IV.1 one would expect the single-blockage phenomenology explored from the periodic boundary case translates to the open boundary case. The lattice can then be found in either of two states: a paused state (bearing similarities with the paused state in periodic boundary conditions), where the lattice is completely filled up to the site of the first paused particle, and a unpaused state, where the system behaves as a standard TASEP. As in the periodic case, the cluster in the paused state consists of paused particles interspersed with $d$ active ones. When the leading paused particle unpauses, it triggers the detachment of $d+1$ particles. This process repeats until the cluster dissolves, after which new particles can re-initiate on the lattice.
Because of the time separation between $\alpha$, $k_{u}$, $k_{p}$ and $\varepsilon$, we can assume that particles quasi-instantaneously fill the lattice up to the position of the first particle becoming paused. Similarly, and for the same reason, particles downstream from the rightmost paused particle would quasi-instantaneously exit the lattice. Reproducing the reasoning followed to obtain Eq. (11), we therefore introduce a two term approach:
$J=\frac{\tau_{0}}{\tau_{p}+\tau_{0}}J_{0}+\frac{\tau_{p}}{\tau_{p}+\tau_{0}}J_{p}\,,$ (18)
@@
where $\tau_{0}$ and $\tau_{p}$ are the average duration of the unpaused state and of the paused state, respectively. The unpaused state behaves as a standard TASEP. Thus, $J_{0}$ is given by the standard TASEP current Eq.(1).
@@
We denote by $x$ the position at which, starting from the unpaused state, a particle transitions to the paused state. The position $x$ (in units of lattice sites) also corresponds to the total number of particles constituting the initial cluster. As a first approximation, we neglect the transient and assume the system enters steady state immediately after exiting the paused state. This leads to a homogeneous particle density, and since all particles have equal probability of pausing, the average position of the first paused particle is $x\right>=L/2$.
Thus the total current of the paused state can be computed as the number of particles in the initial cluster divided by the typical time $\tau_{p}$ of its duration, $J_{p}=L/(2\tau_{p})$. Equation (18) then becomes:
$J=\frac{\tau_{0}}{\tau_{0}+\tau_{p}}J_{0}+\frac{1}{\tau_{0}+\tau_{p}}\frac{L}{2}.$
@@
The first timescale $\tau_{0}$ can be computed by considering that the rate of switching from the unpaused state is proportional to the total number of particles times $k_{p}$, that is
$\tau_{0}=\frac{1}{\rho Lk_{p}}=\frac{\varepsilon}{\alpha Lk_{p}}\,,$
@@
where in the last equivalence we consider a lattice in the LD phase ($\rho=\alpha/\varepsilon$). We then compute the mean lifespan $\tau_{p}$ of the paused state following a numerical approach. This timescale can be smaller or comparable to the time necessary to reach an equilibrium distribution of the paused particle fraction $f_{p}$ given in Eq. (3), and for this reason writing a closed form for $\tau_{p}$ is cumbersome. Thus, after the first pausing event leading to the formation of the extensive cluster, other particles can become paused, slowing down the restoration of the unpaused. Eventually, the paused state will disappear when all particles will successively re-become active and exit the lattice together with the particles that were blocked behind them. The silent assumption behind our reasoning is that active particles which are not trapped in the cluster will exit the lattice without becoming paused again.
@@
To develop our numerical method we first compute the time dependence of $f_{p}$ starting from a situation in which all particles in the cluster are active:
$f_{p}(t)=\frac{k_{p}}{k_{p}+k_{u}}\left(1-e^{-(k_{p}+k_{u})t}\right)\,,$ (19)
@@
which asymptotically reaches the steady-state probability $f_{p}=\frac{k_{p}}{k_{p}+k_{u}}$. At a given time $t$, the mean number $d_{t}$ of particles ‘trapped’ between the first and the second rightmost paused particles can be computed as the mean number of failures in a sequence of independent and identically distributed Bernoulli trials before another particle is found paused (success). This is computed as the average of a negative binomial distribution with rate of success $f_{p}(t)$ (Eq. (19), giving
@@
The average lifetime $\tau(x)$ of a paused particle situated at $x$ is $t_{-}:=1/k_{u}$, and the expected number of particles between two paused particles is given by Eq. (20) computed at $t=t_{-}$. We now have all the elements to numerically compute the typical time to free all particles blocked in the paused state. The number of particles freed after the leading particle becomes active, given after time $t_{1}:=1\times t_{-}$, is given by $d_{1}+1$, where $d_{1}$ is the number of particles in between the first and the second paused particles. If $d_{1}+1$ is greater than $x$, we stop and estimate $\tau(x)=t_{-}$. Otherwise, we compute the next expected distances $d_{i}$ evaluating Eq.(20) at $t_{i}=i\times t_{-}$ until $\sum_{i=1}^{j}(d_{i}+1)>x$. Then, we approximate $\tau(x)$ as $j\times t_{-}$. Averaging $\tau(x)$ over all possible cluster lengths finally gives us the $\tau_{p}$ to be used in Eq. (18).
Figure 9: Compares the current as a function of $k_{p}$ for different system sizes. Here $k_{u}=10^{-3},\alpha=0.1,$ and $\beta=1$. The solution given by Eq. 18 works well for small system sizes and becomes worse for larger lattices.
Our theory is in good agreement with the simulations, particularly for small lattices. In particular, in Fig. 9 we show how our two-state approximation Eq. (18) correctly predicts the particle current $J$ even for small systems ($L=50$) and when the pausing rate $k_{p}$ is varied over several orders of magnitudes.
@@ V Discussions
Intriguingly, in both periodic and open boundary conditions (see Figs. 3-4), this mean-field approximation fails when $k_{p}/\varepsilon$ and $k_{u}/\varepsilon\ll 1$. We provide a few estimates of rates involved in transcription and translation, and point out that biologically relevant regimes are close to, or fall into, this region of poor agreement between theory and simulations. Notably, our model treats particles as occupying a single site, whereas RNAPs and ribosomes each span multiple sites. Accounting for their true footprint would reduce the maximum particle load on the lattice, amplifying finite-size effects and driving both transcription and translation further into the finite-size regime [37]. Corroborated by a systematic exploration of the parameter space (Figs. 5-7), we conclude that severe finite-size effects cause the inadequacy of existing mean-field approaches when $k_{p}/\varepsilon$ and $k_{u}/\varepsilon\ll 1$. In contrast to the standard TASEP, where the mean-field theory performs well even with relatively small lattices (of the order of 100100100 sites), and where the agreement with simulations is independent of hopping and boundary rates, the pTASEP exhibits finite-size effects that depend on parameter values. While the thermodynamic limit remains a reasonable approximation for finite systems, the required system size to achieve this condition is closely tied to the separation of timescales between the particle number (i.e the system size), pausing/unpausing dynamics and particle hopping.
We develop a single-cluster approximation (Section IV.2.1) to better understand the nature of the interplay between this timescale separation and finite-size effects. As the size dependence is related to the disappearance of paused particles in the lattice, our method, which considers the lattice as being in two states (with or without paused particles), captures the relation between current and lattice size. To develop this, in the regime $k_{p}/\varepsilon$ and $k_{u}/\varepsilon\ll 1$ we assume the extreme cases of all particles regrouping in a single cluster, or relaxing to the standard TASEP dynamics when there are no paused particles. This single-cluster approximation reproduces well the simulated current-density relationships when the hypotheses are met (Figs. 5-7). However, the relaxation process (necessary to achieve the steady-state current) being fast compared to the pausing/unpausing kinetics (see Fig. 11), we can generalise our approach to the multi-cluster regime, for which we approximating the current $J_{0}$ in the absence of paused particles by that of the standard TASEP Eq. (1).
@@
Finally, we present the extension of the single-cluster approximation to the open boundary conditions, showing how our two-state hypothesis can be extended to this situation. For instance, our model describes well how the current depends on the pausing rate $k_{p}$ across different orders of magnitudes.